# Continual Learning for Multiple Modalities

### v1

@@ Abstract
Continual learning aims to learn knowledge of tasks observed in sequential time steps while mitigating the forgetting of previously learned knowledge. Existing methods were proposed under the assumption of learning a single modality (e.g., image) over time, which limits their applicability in scenarios involving multiple modalities. In this work, we propose a novel continual learning framework that accommodates multiple modalities (image, video, audio, depth, and text). We train a model to align various modalities with text, leveraging its rich semantic information. However, this increases the risk of forgetting previously learned knowledge, exacerbated by the differing input traits of each task. To alleviate the overwriting of the previous knowledge of modalities, we propose a method for aggregating knowledge within and across modalities. The aggregated knowledge is obtained by assimilating new information through self-regularization within each modality and associating knowledge between modalities by prioritizing contributions from relevant modalities. Furthermore, we propose a strategy that re-aligns the embeddings of modalities to resolve biased alignment between modalities. We evaluate the proposed method in a wide range of continual learning scenarios using multiple datasets with different modalities. Extensive experiments demonstrate that ours outperforms existing methods in the scenarios, regardless of whether the identity of the modality is given.
@@ 1 Introduction
Recently, a vision-language model, pre-trained with a large number of image-text pairs, has drawn significant attention [28] due to its strong generalization performance [18, 29]. Leveraging the general knowledge of the pre-trained model, several algorithms have achieved promising results in image classification [14], semantic segmentation [21], and video action recognition [29], to name a few. However, when deep learning models are exposed to new data, they tend to overwrite previously learned knowledge, leading to forgetting earlier information. This issue is known as catastrophic forgetting [22, 16].
@@
To alleviate this phenomenon, continual learning methods strive to preserve existing knowledge while accommodating new information [16, 1, 30, 33, 11, 17]. To this end, they apply penalties to minimize changes in key parameters that contribute to performance on previous tasks [16, 1] or distill the knowledge of the old model into newer ones [11, 17]. Furthermore, to facilitate recall of previous knowledge, some methods rely on having access to data from old tasks [23, 30, 7, 20]. The collected data is then used jointly with new data to train the model. However, storing raw data can lead to privacy concerns in many applications [34]. Moreover, temporal data (e.g., video and audio) demand more memory resources than non-temporal data [37, 13], posing additional challenges for continual learning.
@@
Recently, continual learning methods that adapt a pre-trained model (e.g., ViT [6] or CLIP [28]) have been proposed [44, 43, 40]. These methods are designed to select prompts from a shared pool for each task. The selected prompts are inserted into a pre-trained model and updated to acquire new knowledge. Despite outperforming other methods by leveraging the general knowledge of the pre-trained model, these approaches validate on a single modality and do not account for data from new modalities. This limitation restricts their applicability in dynamic environments where data from different modalities appear, such as in autonomous driving [26] and robot planning [31]. Furthermore, these approaches are susceptible to forgetting knowledge learned previously when learning multiple modalities because modality gaps between tasks cause prompts trained on different input types to interfere with each other.
@@
In this paper, we propose a novel approach, COMM (COntinual Learning for Multiple Modalities), for continuously learning tasks of different modalities, using a vision-language model. Specifically, we take text-bound data as input, leveraging the rich semantic information provided by text [53]. We process the text data through a language encoder and the other modalities through a modality encoder, enabling the integration of additional modalities without introducing many parameters. However, directly updating the parameters of the encoder across modalities can lead to interference due to differing input traits. To address this issue, we propose a prompt aggregation strategy that associates knowledge within modality through self-regularization, gradually acquiring newly introduced knowledge. It then integrates relevant knowledge between modalities while minimizing the influence of unrelated ones. We introduce a projection head for each modality to assist in aligning embeddings from both encoders, following the practice in [8, 9]. Additionally, to avoid the biased projection of modality data towards text data learned most recently [46], we present a projection head alignment strategy that re-aligns previous modality embeddings with their corresponding text embeddings. Figure 1 illustrates the proposed multimodal continual learning compared to existing unimodal continual learning.
@@
We demonstrate the proposed method by comparing it with existing continual learning methods across tasks involving multiple modalities using classification benchmark datasets: ImageNet-100 [5], UCF-101 [36], SUN-RGBD [35], and ESC-50 [25]. Experimental results show that the proposed method outperforms its competitors across various class-incremental learning scenarios while introducing negligible memory overhead from the prompts. The contributions of our work are four-fold:
@@
  • We self-regulate prompts to consolidate them and selectively prioritize those from relevant modalities, enabling optimal use of prompts within and across modalities.
@@ 2 Related Work
Multimodal learning. Multimodal learning aims to align representations of multiple modalities [51, 14, 29, 19, 45, 10]. They have recently gained attention with the emergence of a model pre-trained on image-text pairs [28]. Several methods [14, 51, 52, 45] have been proposed to adapt the pre-trained model to downstream image classification tasks. Additionally, methods [19, 29] have been proposed to adapt image-text matching models to facilitate video-text matching. They fine-tune pre-trained weights or insert learnable prompts to adapt to downstream tasks. Some recent works [10, 9, 53] have proposed learning multiple modalities (three or more, including image and text). These methods integrate representations extracted using multiple networks, each tailored to a specific modality. Although the methods train modalities simultaneously, training them sequentially presents a greater challenge, as it requires maintaining consistent representations for new modality data while preserving those learned from previously encountered modalities. Additionally, employing a modality-specific network limits scalability as the number of modalities increases. In this work, we present a global encoder for non-text modalities that aligns prompts from relevant modalities, reducing interferences and promoting consensus across different modalities.
@@
Continual learning. The goal of continual learning is to maintain previously learned knowledge while simultaneously acquiring new information. Most existing continual learning methods were proposed based on a single modality (e.g., image). Early works proposed to identify contributing parameters for previous tasks [16, 48, 2, 1]. These methods train the model by penalizing updates to key parameters, helping to alleviate the forgetting of previously learned knowledge. However, they can result in significant loss of knowledge as the number of tasks increases when the semantic domains of the tasks differ [4]. Several studies store [7, 30, 20, 24] or generate [23, 32, 42] some data from old tasks and train the network with a joint set of old and new data. However, privacy or memory budget issues can arise from storing old data [37]. Additionally, these approaches learn a single modality, which limits their applicability in real-world data involving multiple modalities.
@@
Continual learning with a vision-language model. Recent approaches have been proposed using a pre-trained vision-language model [28] to classify unimodal data by matching it with text [50, 45, 38, 40, 27]. Some methods [50, 45] directly update the parameters of a model while regularizing to retain the learned information of the model. The others [38, 40, 27] learn new tasks by combining trainable prompts with the input embeddings in each layer of a pre-trained model, selecting prompts from a predefined pool. However, these methods overwrite previously learned knowledge by updating the parameters of the model (or previously learned prompts), thereby exacerbating forgetting due to interference from different modalities across tasks. We aggregate prompts from modalities relevant to the input by ensuring that knowledge learned from other modalities remains unchanged, alleviating interference between modalities.
@@ 3 Methodology
Figure 2: An illustration of the proposed method for learning multiple modalities ($D^{t}_{\color[rgb]{0,1,1}\textbf{depth}},D^{t}_{\color[rgb]{0,1,0}\textbf{video}},...$) at time step $t$. We introduce a prompt for each modality, which is aggregated with existing prompts, and perform self-regulation to preserve previous knowledge of modalities. To prioritize the knowledge of relevant modalities, we associate the aggregated prompts across modalities. The data and the aggregated prompts are fed into the modality and language encoders to obtain modality and text embeddings, respectively. We also present projection head re-alignment to alleviate biased projections toward the most recently learned text.
@@ 3.1 Problem Definition
[…]
Modality-text alignment. To extract embeddings of the modality data $x^{t}_{m,i}$ and the text data $c^{t}_{text,j}$, we use the modality [^1] We implement the modality encoder using the vision encoder in [28]. and language encoders, $V(\cdot)$ and $L(\cdot)$, in a pre-trained model [28], respectively. This approach helps align embeddings from non-text modalities with those from text, following [9, 53]. To mitigate the increase in memory cost as the number of modalities grows, non-text modalities share a common modality encoder. $x^{t}_{m,i}$ and $c^{t}_{text,j}$ are preprocessed and tokenized as $e^{t}_{m,i}$ and $w^{t}_{text,j}$. By using the tokenized data, we obtain the modality feature $\bar{v}^{t}_{m,i}=V(e^{t}_{m,i})$ and the text embedding $l^{t}_{text,j}=L(w^{t}_{text,j})$, respectively. $\bar{v}^{t}_{m,i}$ is further projected to a common modal-text embedding space as $v^{t}_{m,i}=f_{\alpha^{t}_{m}}(\bar{v}^{t}_{m,i})$, where $f_{\alpha^{t}_{m}}(\cdot)$ is a projection head parameterized by $\alpha^{t}_{m}$. To learn the modality data, we maximize the predictive probability defined as
$p(y^{t}_{i}|x^{t}_{m,i})=\frac{\mbox{exp}(\mbox{sim}(v^{t}_{m,i},l^{t}_{text,y^{t}_{i}})/\tau)}{\sum_{j}\mbox{exp}(\mbox{sim}(v^{t}_{m,i},l^{t}_{text,j})/\tau)},$ (1)
where $l^{t}_{text,y^{t}_{i}}$ denotes the text embedding of the label $y^{t}_{i}$. $\mbox{sim}(\cdot)$ and $\tau$ represent the cosine similarity and the temperature parameter of the pre-trained model, respectively.
@@ 3.2 Framework
The proposed method learns multimodal tasks based on a pre-trained model. However, directly updating the model on different modalities [16] can exacerbate forgetting due to modality gaps and differing input sizes across tasks. To resolve it, we take a strategy of introducing prompts, learnable parameters, to learn new tasks without updating the model [14]. Figure 2 illustrates the proposed continual learning method for multiple modalities.
@@
Specifically, the set of modalities of the task [^2] We assume that a task appears at every time step in continual learning. at time $t$ is $\mathcal{M}^{t}$, where each modality consists of text-paired data. To learn new tasks, we introduce and update a pair of learnable prompts, $P^{t}_{m}$ and $Q^{t}_{m}$, of the modality and language encoders by incorporating them with the input tokens, respectively. Unlike existing prompt-based continual learning [52, 40], which employs a shared prompt pool for unimodal tasks, multimodal continual learning requires a shared pool that can accommodate tasks with diverse modalities. This introduces a challenge in selecting appropriate prompts.
@@
To effectively utilize prompts for each modality and minimize interference with previously learned representations, we propose aggregating prompts within and between modalities in the proposed pipeline. In the first stage, to alleviate significant changes in old knowledge across time within a modality, we combine the prompts learned so far with newly introduced ones by a self-regulating strategy. Note that updating prompts across different modalities can significantly alter the knowledge of the prompts due to gaps in input characteristics, making it more susceptible to forgetting. To address this, in the second stage, we associate knowledge between modalities by leveraging relevant modality connections. We obtain modality and text embeddings using prompts aggregated within and between modalities. The embeddings are aligned by updating $P^{t}_{m}$ and $Q^{t}_{m}$ using the cross-entropy loss,
$\mathcal{L}_{m}=-\sum_{i}y^{t}_{i}\log p(y^{t}_{i}|x^{t}_{m,i}).$ (2)
3.3 Knowledge Aggregation
The use of prompts from a shared prompt pool [44, 40] over time presents two challenges: 1) difficulty in learning new tasks due to the inadequate utilization of existing prompts, and 2) performance degradation as prompts learned for earlier tasks are updated. To address these issues, we propose a knowledge aggregation strategy that minimizes interference for each modality and across modalities over time.
@@ 3.3 Knowledge Aggregation
Aggregation within a modality. To maintain earlier representations and ensure effective prompt utilization within a modality, we aggregate prompts over time and self-regulate them. Specifically, prompts $P^{t}_{m}$ and $Q^{t}_{m}$ are zero-initialized to gradually acquire knowledge of the task. These prompts are then integrated with those from the previous tasks as ¯$\bm{\bar{}}{P}^{t}_{m}=\oplus(\bm{\bar{}}{P}^{t-1}_{m},P^{t}_{m})$ and ¯$\bm{\bar{}}{Q}^{t}_{m}=\oplus(\bm{\bar{}}{Q}^{t-1}_{m},Q^{t}_{m}$), where ¯$\bm{\bar{}}{P}^{1}_{m}=P^{1}_{m}$ and ¯$\bm{\bar{}}{Q}^{1}_{m}=Q^{1}_{m}$. $\oplus$ denotes element-wise summation. Prompts that were not introduced during previous time steps are excluded from the summation. The modality-text embeddings, $V^{t}_{i,\bm{\bar{}}{P}^{t^{\prime}}_{m}}$triangleq $f_{\alpha^{t}_{m}}(V(e^{t}_{m,i};\bm{\bar{}}{P}^{t^{\prime}}_{m}))$ and $L^{t}_{i,\bm{\bar{}}{Q}^{t^{\prime}}_{m}}$triangleq $L(w^{t}_{text,y^{t}_{i}};\bm{\bar{}}{Q}^{t^{\prime}}_{m})$, are aligned by minimizing the loss, $\mathcal{L}_{m}$. However, updating the prompts by the loss does not guarantee the preservation of previous representations. To closely align the integrated knowledge with the previous knowledge, we update prompts with self-regularization as
$\mathcal{L}_{sr}=\sum_{i}||V^{t}_{i,\bm{\bar{}}{P}^{\bm{t-1}}_{m}}-V^{t}_{i,\bm{\bar{}}{P}^{\bm{t}}_{m}}||_{2}+||L^{t}_{i,\bm{\bar{}}{Q}^{\bm{t-1}}_{m}}-L^{t}_{i,\bm{\bar{}}{Q}^{\bm{t}}_{m}}||_{2}.$ (3)
By incorporating the self-regularizations into $\mathcal{L}_{m}$, we derive a representative prompt $\bar{P}^{t}_{m}$ for each modality $m$ that is free from incorrect prompt selection within the modality. Similarly, we aggregate old and new projection heads to form ¯$\bm{\bar{}}{\alpha}^{t}_{m}$= (¯$\oplus(\bm{\bar{}}{\alpha}^{t-1}_{m},\alpha^{t}_{m})$ and apply self-regularization to minimize the distance between outputs extracted using ¯$\bm{\bar{}}{\alpha}^{t}_{m}$ and ¯$\bm{\bar{}}{\alpha}^{t-1}_{m}$ for the input $V(e^{t}_{m,i};\bm{\bar{}}{P}^{t}_{m})$. This approach ensures that both the prompts and projection heads evolve gradually, retaining previously learned knowledge within the modality.
@@
Aggregation between modalities. Prompts trained across modalities are more sensitive to inappropriate usage than those trained for a single modality. This necessitates the identification of relevant modalities from the pool observed thus far, denoted as $\mathcal{M}^{1:t}$. To do so, we introduce a function $g_{\beta^{t}}(\cdot)$: $\mathbb{R}^{d_{i}}\rightarrow\mathbb{R}^{|\mathcal{M}^{1:t}|}$, parameterized with $\beta^{t}$, to identify relevant modalities by examining the features extracted from the modality encoder, $V(\cdot)$. The function $g_{\beta^{t}}(\cdot)$ assigns probabilities to the prompts, representing their relevance and facilitating the integration of prompts from the modalities. Due to the inaccessibility of old data, we collect features for each modality, which are sampled [49] from a normal distribution defined by the mean and covariance of the previously observed features, to update $g_{\beta^{t}}(\cdot)$
$\mathcal{L}_{agg}=\sum_{m^{\prime}\in\mathcal{M}^{1:t}}\sum_{i=1}^{\bar{n}_{m^{\prime}}}~{}-\log g_{\beta^{t}}(m^{\prime}|\bar{u}_{m^{\prime},i}),$ (4)
@@
where $\bar{u}_{m^{\prime},i}$ is a collected sample and $\bar{n}_{m^{\prime}}$ denotes the number of samples. By minimizing this loss, $g_{\beta^{t}}$ yields relevance of each modality $m^{\prime}\in\mathcal{M}^{1:t}$ in relation to $m$. Then we aggregate the prompts across modalities with the relevance $g_{\beta^{t}}$ as:
^$\bm{\hat{}}{P}^{t}_{m}=\sum_{m^{\prime}\in\mathcal{M}^{1:t}}g_{\beta^{t}}(m^{\prime}|V(e^{t}_{m,i}))\cdot\bm{\bar{}}{P}^{t}_{m^{\prime}},$ (5)

### v2

@@ Abstract
Continual learning aims to learn knowledge of tasks observed in sequential time steps while mitigating the forgetting of previously learned knowledge. Existing methods were designed to learn a single modality (e.g., image) over time, which limits their applicability in scenarios involving multiple modalities. In this work, we propose a novel continual learning framework that accommodates multiple modalities (image, video, audio, depth, and text). We train a model to align various modalities with text, leveraging its rich semantic information. However, this increases the risk of forgetting previously learned knowledge, exacerbated by the differing input traits across tasks. To alleviate the overwriting of previous knowledge of modalities, we propose a framework that consolidates intra-modal knowledge while incorporating relevant inter-modal information. This is achieved by self‑regulating shifts in learned representations to gradually integrating novel knowledge into the information retained across modalities. Simultaneously, it mitigates inter-modal interference by selectively integrating knowledge from previously encountered modalities based on their mutual relevance. Furthermore, we introduce a strategy to re-align modality embeddings, effectively addressing biased alignment between modalities. We evaluate the proposed method in a wide range of continual learning scenarios using multiple datasets with different modalities. Extensive experiments demonstrate that ours outperforms existing methods in the scenarios, regardless of whether the identity of the modality is given.
@@ 1 Introduction
Recently, a vision-language model, pre-trained with a large number of image-text pairs, has drawn significant attention (Radford et al. 2021) due to its strong generalization performance (Lin et al. 2023; Rasheed et al. 2023). Leveraging the general knowledge of the pre-trained model, several algorithms have achieved promising results in image classification (Khattak et al. 2023), semantic segmentation (Lüddecke and Ecker 2022), and video action recognition (Rasheed et al. 2023), to name a few. However, when deep learning models are exposed to new data, they tend to overwrite previously learned knowledge, leading to forgetting earlier information. This issue is known as catastrophic forgetting (McCloskey and Cohen 1989; Kirkpatrick et al. 2017).
@@
To alleviate this phenomenon, continual learning methods strive to preserve existing knowledge while accommodating new information (Kirkpatrick et al. 2017; Aljundi et al. 2018; Rebuffi et al. 2017; Smith et al. 2023a; Hou et al. 2018; Li and Hoiem 2016). To this end, they apply penalties to minimize changes in key parameters that contribute to performance on previous tasks (Kirkpatrick et al. 2017; Aljundi et al. 2018) or distill the knowledge of the old model into newer ones (Hou et al. 2018; Li and Hoiem 2016). Furthermore, to facilitate recall of previous knowledge, some methods rely on having access to data from old tasks (Ostapenko et al. 2019; Rebuffi et al. 2017; Douillard et al. 2022; Liu, Schiele, and Sun 2021). The collected data is then used jointly with new data to train the model. However, storing raw data can lead to privacy concerns in many applications (Smith et al. 2023b). Moreover, temporal data (e.g., video and audio) demand more memory resources than non-temporal data (Villa et al. 2022; Jin et al. 2023), posing additional challenges for continual learning.
@@
Recently, continual learning methods that adapt a pre-trained model (e.g., ViT (Dosovitskiy et al. 2020) or CLIP (Radford et al. 2021)) have been proposed (Wang et al. 2022, 2023; He, Duan, and Zhu 2025). These methods are designed to select prompts from a shared pool for each task. The selected prompts are inserted into a pre-trained model and updated to acquire new knowledge. Despite outperforming other methods by leveraging the rich knowledge of the pre-trained model, these approaches validate on a single modality and do not account for data from new modalities. This limitation restricts their applicability in dynamic environments where data from different modalities appear, such as in autonomous driving (Prakash, Chitta, and Geiger 2021) and robot planning (Huang et al. 2025). More importantly, these approaches are susceptible to forgetting knowledge learned previously when learning multiple modalities because modality gaps between tasks cause prompts trained on different input types to interfere with each other.
@@
In this paper, we propose a novel approach, COMM (COntinual Learning for Multiple Modalities), for continuously learning tasks of different modalities, using a pre-trained vision-language model. Specifically, we take text-bound data as input, leveraging the rich semantic information provided by text (Zhu et al. 2023). We process the text data through a language encoder and the other modalities through a modality encoder, enabling the integration of additional modalities without introducing many parameters. To leverage rich pre-trained knowledge, we introduce prompts to accommodate new knowledge without updating the backbone parameters. These prompts are accumulated over time for each modality through self-regularization, which preserves representational consistency across time and supports stable integration of intra-modality knowledge. To reduce interference across modalities, we promote modality-aware adaptation by aggregating prompts from previously encountered modalities, suppressing the influence of unrelated ones. Additionally, to prevent biased alignment of recently learned text (Wu et al. 2019), we re-align earlier modality representations to restore semantic consistency with their textual counterparts. Figure 1 illustrates the proposed multimodal continual learning compared to existing unimodal continual learning.
@@
We demonstrate our method by comparing it with existing continual learning methods across tasks involving multiple modalities using classification benchmark datasets: ImageNet-100 (Deng et al. 2009), UCF-101 (Soomro, Zamir, and Shah 2012), SUN-RGBD (Song, Lichtenberg, and Xiao 2015), and ESC-50 (Piczak 2015). Experimental results show that the proposed method outperforms its competitors across various class-incremental learning scenarios while introducing negligible memory overhead from the prompts. The contributions of our work are four-fold:
@@
  • Our re-alignment strategy restores distorted semantic consistency between previously learned modalities and text during adaptation to new tasks.
@@ 2 Related Work
Multimodal learning aims to construct unified representations from diverse data types, with significant progress driven by pre-trained models such as CLIP (Radford et al. 2021). These models are commonly adapted to downstream tasks, including image classification (Khattak et al. 2023; Zhou et al. 2022b; Wortsman et al. 2022) and video-text matching (Lin et al. 2022; Rasheed et al. 2023), via fine-tuning pre-trained weights or inserting learnable prompts. Recent studies (Girdhar et al. 2023; Zhu et al. 2023; Zhou et al. 2025) have extended this framework to encompass three or more modalities, including image and text. While these approaches train modalities jointly, sequential training poses a greater challenge, as it must preserve representations of previously learned modalities while integrating new ones. Moreover, existing methods often rely on modality-specific networks (Girdhar et al. 2023; Zhu et al. 2023), limiting scalability as the number of modalities grows. In this work, we present a global encoder for non-text modalities that aligns prompts across modalities, mitigating interference and enhancing representational consensus.
@@
Continual learning aims to maintain previously learned knowledge while simultaneously acquiring new information. Existing continual learning methods have primarily focused on a single modality (e.g., image). Early works proposed to identify contributing parameters for previous tasks (Kirkpatrick et al. 2017; Friedemann, Poole, and Ganguli 2017; Aljundi et al. 2018). These methods train the model by penalizing updates to key parameters, which alleviates forgetting of previously learned knowledge. However, they can cause significant loss of knowledge as the number of tasks with different semantic domains increases (De Lange et al. 2021). Several studies store (Douillard et al. 2022; Liu, Schiele, and Sun 2021; Park, Kang, and Han 2021) or generate (Ostapenko et al. 2019; Shin et al. 2017; Wang et al. 2019) some data from old tasks and train the network with a joint set of old and new data. However, privacy or memory budget issues can arise from storing old data (Villa et al. 2022). Importantly, the unimodal focus of these approaches limits their applicability in real-world settings where data arrives as a continuous stream from multiple modalities.
Figure 2: An illustration of the proposed method for learning multiple modalities at time step $t$. The model receives modality data paired with the corresponding text. First, new prompts are introduced and aggregated with previously accumulated ones via self-regularization to preserve intra-modal knowledge. Then, intra-modal prompts are integrated into final prompts using modality-aware relevance weights, which suppress cross-modal interference. The resulting prompts are passed into the modality encoder to obtain aligned features. To prevent biased projections toward recently learned texts, projection heads are re-aligned to restore semantic consistency with previously learned alignments.
@@
Continual learning with a vision-language model (Radford et al. 2021) has recently been explored, with methods adapting to new unimodal data by matching it with corresponding text descriptions (Wortsman et al. 2022; Villa et al. 2023; Wang et al. 2023; Wu et al. 2025). Some methods (Zheng et al. 2023; Wortsman et al. 2022) directly update the parameters of a model while regularizing to retain its learned information. The others (Villa et al. 2023; Wang et al. 2023; Qiao et al. 2024) learn new tasks by combining trainable prompts with the input embeddings in each layer of a pre-trained model, selecting prompts from a predefined pool. However, these methods often overwrite previously learned knowledge by updating model parameters or prompts, exacerbating forgetting due to interference from different modalities across tasks. In contrast, we aggregate prompts from modalities relevant to the input, ensuring that knowledge learned from other modalities remains unchanged, alleviating interference between modalities.
@@ 3.1 Problem Definition
We present a new problem of learning multiple modalities over a sequence of tasks, where at each time step $t$, a subset of modalities, $\mathcal{M}^{t}\subseteq\mathcal{M}=\{$image, video, depth, audio,},...\}, is presented. Specifically, at time step $t$, we learn knowledge of each modality $m\in\mathcal{M}^{t}$ by aligning it with the corresponding class name (represented by text) in a joint embedding space. The goal of multimodal continual learning is to maintain previously learned knowledge (modality-text alignments) while integrating new knowledge from the current task. At time step $t$, the task [^1] We assume that a task appears at every time step. is defined by a set of modality-text pairs $T^{t}=\{D^{t}_{m},C^{t}_{text}\}_{m\in\mathcal{M}^{t}}$. For each modality $m$, $D^{t}_{m}$ comprises data points and their labels $\{x^{t}_{m,i},y^{t}_{i}\}_{i=1}^{n^{t}_{m}}$, and $C^{t}_{text}$ provides the corresponding textual class names $\{x^{t}_{\text{text},i}\}_{i=1}^{n^{t}_{m}}$.
Modality-text alignment. To extract embeddings of the modality data $x^{t}_{m,i}$ and the text data $x^{t}_{\text{text},i}$, we use the modality [^2] We implement the modality encoder using the vision encoder in (Radford et al. 2021). and language encoders, $V(\cdot)$ and $L(\cdot)$, in a pre-trained model (Radford et al. 2021), respectively. This approach helps align embeddings from non-text modalities with those from text (Girdhar et al. 2023; Zhu et al. 2023). To mitigate the increase in memory cost as the number of modalities grows, non-text modalities share a common modality encoder. We obtain the modality feature $\tilde{v}^{t}_{m,i}=V(x^{t}_{m,i})$ and the text embedding $l^{t}_{\text{text},i}=L(x^{t}_{\text{text},i})$, respectively. $\tilde{v}^{t}_{m,i}$ is further projected to a common modal-text embedding space as $v^{t}_{m,i}=f_{\alpha^{t}_{m}}(\tilde{v}^{t}_{m,i})$, where $f_{\alpha^{t}_{m}}(\cdot)$ is a projection head parameterized by $\alpha^{t}_{m}$. To learn the modality data, we maximize the predictive probability defined as
$p(y^{t}_{i}|x^{t}_{m,i})=\frac{\mbox{exp}(\mbox{sim}(v^{t}_{m,i},l^{t}_{\text{text},i})/\tau)}{\sum_{j}\mbox{exp}(\mbox{sim}(v^{t}_{m,i},l^{t}_{\text{text},j})/\tau)},$ (1)
where $\mbox{sim}(\cdot)$ and $\tau$ represent the cosine similarity and the temperature parameter, respectively.
@@ 3.2 Framework
The proposed method learns multimodal tasks based on a pre-trained model. However, directly updating the model on different modalities can exacerbate forgetting due to interference caused by modality gaps; representational discrepancies and conflicting updates from informational imbalance (Peng et al. 2022). To resolve it, we take a strategy of introducing prompts to learn new tasks without updating the model itself (Khattak et al. 2023). Figure 2 illustrates our continual learning method for multiple modalities.
@@
Specifically, the set of modalities of the task at time $t$ is $\mathcal{M}^{t}$, where each modality consists of text-paired data. To learn new tasks, we introduce and update a pair of learnable prompts, $P^{t}_{m}$ and $Q^{t}_{m}$, of the modality and language encoders by incorporating them with the input, respectively. Unlike existing prompt-based continual learning (Zhou et al. 2022b; Wang et al. 2023), which employs a prompt pool for unimodal tasks, multimodal continual learning requires a shared pool capable of accommodating tasks with diverse modalities. This introduces a challenge in selecting appropriate prompts.
@@
To effectively utilize prompts for each modality and minimize interference with previously learned representations, we propose aggregating prompts intra and inter modalities in the proposed pipeline. In the first stage, to alleviate significant changes in old knowledge across time within a modality, we combine the prompts learned so far with newly introduced ones by a self-regulating strategy. The update of prompts across modalities can lead to knowledge degradation due to input discrepancies; thus, the second stage mitigates forgetting by aligning knowledge through modality-aware associations. We obtain modality and text embeddings using prompts aggregated intra and inter modalities. The embeddings are aligned by updating $P^{t}_{m}$ and $Q^{t}_{m}$ using the cross-entropy loss, $\mathcal{L}_{\text{task}}=-\sum_{i}y^{t}_{i}\log p(y^{t}_{i}|x^{t}_{m,i}).$
3.3 Continual Learning for Multiple Modalities
@@
Intra-modal aggregation via self-regularization. To maintain earlier representations and ensure effective prompt utilization within each modality, we aggregate prompts over time and self-regulate them. Specifically, prompts $P^{t}_{m}$ and $Q^{t}_{m}$ are zero-initialized to gradually acquire knowledge of the task. These prompts are then integrated with those from the previous tasks as $\bm{\bar{}}{P}^{t}_{m}=\oplus(\bm{\bar{}}{P}^{t-1}_{m},P^{t}_{m})$ and $\bm{\bar{}}{Q}^{t}_{m}=\oplus(\bm{\bar{}}{Q}^{t-1}_{m},Q^{t}_{m}$), where $\bm{\bar{}}{P}^{1}_{m}=P^{1}_{m}$ and $\bm{\bar{}}{Q}^{1}_{m}=Q^{1}_{m}$. $\oplus$ denotes element-wise summation, and prompts absent in earlier time steps are excluded from this operation. The aggregated prompts are used to extract modality-text embeddings, $V_{\bm{\bar{}}{P}^{t}_{m}}(x^{t}_{m,i})$ and $L_{\bm{\bar{}}{Q}^{t}_{m}}(x^{t}_{\text{text},i})$, which are aligned by minimizing the loss, $\mathcal{L}_{\text{task}}$. However, optimizing solely for the task loss may fail to adequately preserve prior knowledge, leading to significant shifts in the accumulated representations.
To address this, we introduce a self-regularization loss, $\mathcal{L}_{\text{self}}$, applied to both the prompts and the projection heads, enforcing consistency with previously accumulated knowledge. Projection head parameters are similarly aggregated over time: $\bm{\bar{}}{\alpha}^{t}_{m}$= $\oplus(\bm{\bar{}}{\alpha}^{t-1}_{m},\alpha^{t}_{m})$. The complete self-regularization loss is then defined as
$\mathcal{L}_{\text{self}}=\sum_{i}\langle V_{\bm{\bar{}}{P}^{t}_{m}},x^{t}_{m,i}\rangle+\langle L_{\bm{\bar{}}{Q}^{t}_{m}},x^{t}_{\text{text},i}\rangle+\langle f_{\bm{\bar{}}{\alpha}^{t}_{m}},\bar{v}_{m,i}^{t}\rangle,$ (2)
where $\langle V_{A_{m}^{t}},b\rangle=\|V_{A_{m}^{t}}(b)-V_{A_{m}^{t-1}}(b)\|_{2}$ denotes the $\ell_{2}$ distance between the outputs of $V_{A_{m}}(\cdot)$ given an input $b$ using parameters at the current and previous time steps, and $\bar{v}_{m,i}^{t}=V_{\bm{\bar{}}{P}^{t}_{m}}(x^{t}_{m,i})$. This loss penalizes output changes of the prompts and projection head by comparing their current outputs with those obtained using the previously aggregated parameters. Incorporating $\mathcal{L}_{\text{self}}$ into $\mathcal{L}_{\text{task}}$ encourages stable knowledge integration into the parameters $\bm{\bar{}}{P}^{t}_{m},\bm{\bar{}}{Q}^{t}_{m}$, and $\bm{\bar{}}{\alpha}^{t}_{m}$. This strategy enables gradual adaptation of prompts and projection heads while it preserves prior knowledge and prevents interference from incorrect prompt selection.
@@
Inter-modal integration guided by relevance. Prompts trained across modalities are more vulnerable to interference when applied in unrelated contexts than those trained for a single modality. This necessitates the identification of relevant modalities from the pool observed thus far, denoted as $\mathcal{M}^{1:t}$. To this end, we introduce a function $g_{\beta^{t}}(\cdot)$: $\mathbb{R}^{d_{i}}\rightarrow\mathbb{R}^{|\mathcal{M}^{1:t}|}$, parameterized with $\beta^{t}$, to identify relevant modalities by examining the features extracted from the modality encoder, $V(\cdot)$. The function assigns probabilities to the prompts, representing their relevance and facilitating the integration of prompts from the modalities $\{\bm{\bar{}}{P}^{t}_{m^{\prime}}\}_{m^{\prime}\in\mathcal{M}^{1:t}}$. Due to the inaccessibility of old data, we collect features for each modality, which are sampled from a normal distribution defined by the mean and covariance of the previously observed features (Zhang et al. 2023). The function $g_{\beta^{t}}(\cdot)$ is trained by minimizing
$\mathcal{L}_{\text{cross}}=\sum_{m^{\prime}\in\mathcal{M}^{1:t}}\sum_{i=1}^{\bar{n}_{m^{\prime}}}~-\log g_{\beta^{t}}(m^{\prime}|\bar{u}_{m^{\prime},i}),$ (3)
where $\bar{u}_{m^{\prime},i}$ is a collected sample and $\bar{n}_{m^{\prime}}$ denotes the number of samples.
@@
By minimizing this loss, $g_{\beta^{t}}$ yields relevance of each modality $m^{\prime}\in\mathcal{M}^{1:t}$ in relation to $m$. The prompts are then aggregated across modalities using the relevance $g_{\beta^{t}}$ as weights:
$\bm{\hat{}}{P}^{t}_{m}=\sum_{m^{\prime}\in\mathcal{M}^{1:t}}g_{\beta^{t}}(m^{\prime}|V(x^{t}_{m,i}))\cdot\bm{\bar{}}{P}^{t}_{m^{\prime}},$ (4)